OpenStudy (anonymous):

hello. I need to solve mn= 3m-2 please?

OpenStudy (anonymous):

What variable are you solving for? (:

OpenStudy (anonymous):

oh, for m

OpenStudy (anonymous):

help, please?

OpenStudy (anonymous):

Okay hold on sorry my parents are talking to me.

OpenStudy (anonymous):

no problem sorry

OpenStudy (anonymous):

you're really not supposed to divide by a variable but i think thats the only way to do this.

OpenStudy (anonymous):

if no solution is an option that would be your answer because even if you divide by n first then divide by m your m's will cancel to out to a 1. i think.

OpenStudy (anonymous):

but if its for n, then your m's cancel out and you get a 1

OpenStudy (anonymous):

Ya thats what i was thinking because she said solve for m...

OpenStudy (anonymous):

\[mn = 3m-2 \iff m(3-n) = 2 \iff m = \frac{2}{3-n}\]

OpenStudy (anonymous):

m=-2/(n-3)

OpenStudy (anonymous):

\[\text{Note } n \not= 3 \]

OpenStudy (anonymous):

M = 2

OpenStudy (anonymous):

ok, thank you all so much!

OpenStudy (anonymous):

Welcome (: haha i just gave you the answer.. He explained it (:

OpenStudy (anonymous):

(M = 2 isn't the answer)

OpenStudy (anonymous):

It isnt!?!?!? :( Of course :P